Adult memberships cost $10. Students $5. The students sold 50 memberships for a total of $400. How many members are adults and how many are students?

Answer by mananth(16946)

x Adult memberships
y Child memberships
Total memberships
1 x + 1 y = 50 .............1
Cost of total memberships
10 x + 5 y = 400 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 1
-5 x -5 y = -250
10 x + 5 y = 400
Add the two equations
5 x = 150
/ 5
x = 30.00
plug value of x in (1)
1 x + 1 y = 50
30 + y = 50
y = 50 -30
y = 20
y = 20.00
30 Adult memberships
20 Child memberships
m.ananth@hotmail.ca